Suppose a major impact blasted out rock from a 140-km-diameter crater that was, on average, 1 km deep. If all that material became dust that settled out evenly over the entire surface of the Earth, how thick a layer would it form?

The layer of dust would be about cm thick.
Answer #1

diameter of crater = 140 km

radius of crater (r)= 140/2 = 70 km = 7 x 104 m

deepness of crater (d)= 1 km = 103 m

now, volume of dust (V) = π r2 d =π x (7 x 104)2 x 103 = π x 49 x 108 x 103 =  π x 49 x 1011 m3

now radius of earth (R) = 6,371 km = 6.371 x 106 m

let the thickness of dust layer = t

now , 4 π R2 t = V

or ,  4 x π x (6.371 x 106 )2 x t = π x 49 x 1011

or, 4 x π x 39.9 x 1012 x t = π x 49 x 1011

or, t =  (49 x 1011 ) / (4 x 39.9 x 1012) = 3.07 x 10-2 m = 3.07 cm

hence, The layer of dust would be about 3.07 cm thick.